On January 1, 2017, Sheffield Corp. issued ten-year bonds with a face amount of $5800000 and a stated interest rate of 8% payable annually on January 1. The bonds were priced to yield 10%. Present value factors are as follows:

Present value of 1 for 10 periods At 8% 0.463.    At 10%. 0.386
Present value of an ordinary annuity of 1 for 10 periods. At 8%  6.710.   At 10%.   6.145

The total issue price of the bonds was:
Answer #1
Bond characterstics Amount
1-a) Principal 5,800,000
interest            464,000
Market interest rate 10%
periods to maturity 10
issue price 5,090,080 answer
Calculation of bond issue price
Where
i= 10.00%
t= 10
principal * PV of $1 at 10% for 10 yrs =
5,800,000 * 0.386        = 2238800
interest * PV of ordinary annuity at 10%=
464000 * 6.145 = 2851280
bond issue price 5090080